Lulabalula:

Eight friends ate StartFraction 5 Over 8 EndFraction of a bag of chips. What fraction of the bag did each person eat, assuming that they each ate the same amount?

Lulabalula:

@supie help plz

ramen:

Can you upload the screenshot plz? It would make this easier to read

Lulabalula:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @ramen Can you upload the screenshot plz? It would make this easier to read \(\color{#0cbb34}{\text{End of Quote}}\)

Cupcake123456:

who do u think comes first dividing or multiplication ?

Lulabalula:

multiplying

Cupcake123456:

because we know 8 friends ate (5/8)TH fraction of a bag of chips. And Each one ate the same fraction.

Cupcake123456:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Lulabalula multiplying \(\color{#0cbb34}{\text{End of Quote}}\) no

Lulabalula:

so dividing comes first

jhonyy9:

8 friends ate 5/8 bag of a bag of chips 1 pers. eat x x = (5/8) /8 = (5/8) *(1/8) = 5/64

xxDeppressionxx:

All of us do not need to crowed this post

ramen:

I agree with Johnny

Cupcake123456:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 8 friends ate 5/8 bag of a bag of chips 1 pers. eat x x = (5/8) /8 = (5/8) *(1/8) = 5/64 \(\color{#0cbb34}{\text{End of Quote}}\) you should let him do it on his own!! thats what i was doing step by step makes him understand more

Lulabalula:

im a girl-

Cupcake123456:

my bad *her

Lulabalula:

thanks anyway